### v1

# Enabling decision support over confidential data

Abstract
Enabling automated decision-making processes by leveraging data-driven analysis is a core goal of Decision Support Systems (DSSs). In multi-party scenarios where decisions rely on distributed and sensitive data, though, ensuring confidentiality, verifiability, transparency, integrity, and consistency at once remains an open challenge for DSSs. To tackle this multi-faceted problem, we propose the Secure Platform for Automated decision Rules via Trusted Applications (SPARTA) approach. By leveraging Trusted Execution Environments (TEEs) at its core, SPARTA ensures that the decision logic and the data remain protected. To guarantee transparency and consistency of the decision process, SPARTA encodes decision rules into verifiable software objects deployed within TEEs. To maintain the confidentiality of the outcomes while keeping the information integrity, SPARTA employs cryptography techniques on notarized data based on user-definable access policies. Based on experiments conducted on public benchmarks and synthetic data, we find our approach to be practically applicable and scalable.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23-27 February 2026, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2026.[23||24]xxxx www.ndss-symposium.org
I Introduction
Decision Support System (DSS) is an information system designed to help decision-makers solve complex problems and optimize decision processes. DSS enhances decision-making by providing data-driven insights, analytical tools, and structured problem-solving support. These systems facilitate efficient decision-making by leveraging IT-based solutions to process large volumes of information, evaluate multiple scenarios, and improve the overall effectiveness of strategic and operational decisions [1, 2]. They play a pivotal role in several fields including healthcare [3, 4], finance [5], business [6], and supply-chain [7] due to their ability to process large volumes of data, generate insights, and enhance decision accuracy. By efficiently evaluating various conditions and scenarios, these systems contribute to improved time and resource management while ensuring robust performance assessments. Ultimately, DSS enhances decision-making by offering reliable, adaptable, and user-centric solutions, even when problem parameters are not well-defined [8]. In multi-party processes, the decision-making support spans across multiple sources of information. In these cases, it is not easy to establish a trusted exchange of data and a common understanding of both the information and the decision logic [9, 10]. To be applicable in sensitive domains such as healthcare or finance, decision support systems must ensure the confidentiality and traceability of data, support user-definable decision logic, and guarantee system security and availability. Classical decision support systems that cater for customization support languages like Decision Model and Notation (DMN) [11], but are typically designed as modules of centralized information systems that do not guarantee privacy or traceability. Custom privacy-oriented approaches have been proposed to safeguard confidentiality and security but cannot provide means for user-definable decision logic or availability [12, 13] Blockchain-based solutions allow for customization and security and availability, but not for confidentiality [14].
In this paper, we present Secure Platform for Automated decision Rules via Trusted Applications (SPARTA), a novel solution that leverages Trusted Execution Environment (TEE) [15] to securely execute automatic decision support systems in multi-party processes. Our TEE-based solution enables automated decision-making support securely and privately while maintaining trust, transparency, and auditability.
To summarize, our main contributions are as follows:
  1. 1.
    We present a novel TEE-based architecture that securely executes decision support processes in multi-party settings, ensuring code integrity, confidential decision-making, and secure data exchange without exposing sensitive information.
  2. 2.
    Our solution handles access control policies that enforce fine-grained, verifiable authorization over customizable decision functions, ensuring that only authorized participants can trigger decision making on confidential data.
  3. 3.
    We provide an open-source implementation of our solution, which we use to measure its performance, evaluate its scalability, and assess the overhead introduced by our secure platform with public, cross-domain benchmarks.
In the remainder of the paper, Sect. II outlines the core concepts we used to devise our solution. Sect. III presents a running example that is a reference throughout the paper and outlines the problem we tackle along with the requirements it poses. In Sect. IV, we describe our approach in detail. Sect. V provides a formal security analysis of our solution. In Sect. VI, we present our proof-of-concept implementation and the results of the experiments we conducted therewith. In Sect. VII, we discuss the implications and limitations of our work. Sect. VIII reviews related work. Finally, Sect. IX concludes the paper and outlines directions for future work. Following the open science principles, the source code and experiment data can be found at: anonymous.4open.science/r/SPARTA.
II Background
Confidential computing is at the core of our approach, supplemented by decision logic and access control, and distributed notarization and storage. Here, we outline the core principles, technologies, and languages underlying these notions.
Confidential computing
Confidential computing represents a paradigm wherein sensitive data undergoes processing within a protected processing unit. Within this unit, the data and the processes treating those data remain concealed from any external entities [16]. Confidential computing underpins the Trusted Execution Environments (TEEs) [15]. A TEE is an environment that provides a level of assurance of the following three properties: (i) data confidentiality, namely the fact that unauthorized entities cannot access or observe data while it resides within the TEE and is being processed, (ii) data integrity, namely the fact that unauthorized entities cannot modify, delete, or insert data while it remains within the TEE and is undergoing processing and (iii) code integrity, namely that unauthorized entities cannot manipulate, delete, or introduce alterations to the code executed within the TEE [17, 18, 19]. Our work leverages Intel Software Guard Extensions (SGX) [20, 21, 22, 23], a collection of security-related instructions embedded within contemporary Intel Central Processing Unit (CPU)s. SGX facilitates the secure execution of code (deployed in the form of so-called Trusted Applications, or Trusted Apps for short) and data within enclaves, namely hardware-protected environments. The three primary security features of enclaves are isolation, attestation, and sealing. Isolation involves the enclave’s placement within a hardware-secured memory enclave page cache, specifically designated for housing protected code and data. However, TEE-allocated memory is limited. Exceeding these limits may entail stops that can severely affect availability (the service is interrupted) or data secrecy (encrypted data can be swapped to outer memory regions). Additionally, SGX enables remote attestation to allow external entities to authenticate a target enclave. During remote attestation, the external entity communicates with the attestation service to validate the integrity of the information provided by the target enclave. This process entails comparing enclave information with the trusted configuration within the target enclave, ensuring the integrity of code and data loading during the initial enclave establishment. Moreover, SGX employs sealing mechanisms to safeguard enclave contents stored beyond the enclave’s boundaries. These mechanisms entail encrypting enclave contents using a concealed sealing key, derivable from either the measurement hash or the author identity of the enclave. Both sealing mechanisms perform Advanced Encryption Standard - Galois/Counter Mode (AES-GCM) encryption. This approach guarantees exclusive access and decryption privileges to the sole corresponding enclave.
Decision logic and access control
One of the most well-known languages to model and formalize decision logic in business applications is Decision Model and Notation (DMN[24, 25, 26], a standard by the Object Management Group (OMG) 111omg.org/spec/DMN, accessed 2025-06-17 for modeling and formalizing decision logic in multi-party processes and business applications. A DMN, at its core, is a decision table that captures critical knowledge and automates decision logic [27]. A DMN table consists of columns representing the inputs and outputs of a decision, and rows denoting rules (the decision logic). Rules are expressed using Friendly Enough Expression Language (FEEL), a language defined within the DMN standard. Each row has an identifier, a condition for each input column, and one specific value for each output column. A rule is considered applicable when the input values satisfy all associated conditions; in that case, the corresponding output values are produced [27, 26]. To express authorization policies over the usage of resources, the Abbreviated Language for Authorization (ALFA) 222alfa.guide/alfa-authorization-language/, accessed 2025-06-17 language provides a concise, human-readable format. Built on top of the eXtensible Access Control Markup Language (XACML),333docs.oasis-open.org/xacml, accessed 2025-06-17 ALFA offers a simplified syntax for defining fine-grained, attribute-based access control (ABAC) policies based on users, actions, resources, and context.
Distributed notarization and storage
A blockchain is a Distributed Ledger Technology (DLT) wherein transactions are sorted, organized into blocks, and interconnected to form a chain. Blockchains possess resistance against tampering, achieved through cryptographic methods. Public blockchains like Ethereum [28] and Algorand [29] permit the deployment and execution of programs named smart contracts [30, 31]. To mitigate the expenses associated with the invocation of smart contracts, external Peer-to-peer (P2P) systems are frequently employed for storing large volumes of data [32]. Among the facilitating technologies is the InterPlanetary File System (IPFS) [33],444 ipfs.tech, accessed 2025-07-30 a distributed system designed for file storage and access, leveraging a Distributed Hash Table (DHT) to distribute files across multiple nodes. Utilizing content-addressing, IPFS assigns a unique identifier to each file within the network. When a file is uploaded to IPFS, it generates a content-addressed resource locator based on its hash (e.g., QmbPg[…]shTg9). Changing the file changes the hash and hence makes the locator invalid. IPFS supports mutable pointers that allow content updates without changing the shared reference. Specifically, InterPlanetary Name System (IPNS) enables the creation of mutable pointers, known as IPNS names (e.g., PatientsWAggrKey: QmbPg[…]shTg9). These names point to the content-addressed locators, and can be updated with newer locators while retaining the verifiability of content addressing.
III Problem illustration, running example, and requirements
Table I: An excerpt of the data fields provided to the system by the process participants of Sect. III
Central medical hub Vaccination center Patient Carrier
Field Type Example Field Type Example Field Type Example Field Type Example
Stock ID Integer 10523 Hub Name String Ayala PLC Full Name String Andre Mcdaniel Carrier Name String PrimeWay Haulage
Vaccine Type String mRNA Location String Berlin, GE Date of Birth Date-time 1976-08-23 Coverage Area Float 2020.7 [km]
Received Date Date-time 2025-07-15 Current Vaccine Stock Level Float 265.0 Pre-existing Conditions String Asthma Number of Trucks Integer 13
Expiry Date Date-time 2026-01-15 Population Served Integer 182252 Current Medications String Albuterol Number of Planes Integer 2
Quantity Available Integer 2400 Max Storage Capacity Integer 2000 Family Medical History String Heart Disease Refrigeration Capacity Boolean True
Storage Temperature Float -20.5 [°C] Vaccination Progress Integer 686 [per day] Consent to Vaccination Boolean True Temperature Monitoring Systems Boolean False
To illustrate the problem we address, we introduce a real-world scenario from the healthcare domain: a vaccine distribution campaign inspired by [34, 35] and linked to a research project funded by the European NextGenerationEU. [^5] To comply with the double-blindness requirement of the conference, we will reveal the project only upon acceptance. We use this scenario as a running example and motivating case study throughout the paper.
In our scenario, a central medical hub (henceforth, medical hub for short) receives vaccine supplies and handles their dispatch in a federal State. The downstream process involves other participants, namely patients who get an injection, vaccination centers who administer the injections, and several specialized carriers delivering the vaccines on premise. It covers two stages.
In the first phase, the involved participants act as Data Providers by contributing data into the system. We will refer to it as the data provision stage. In our scenario, e.g., the medical hub reports received vaccine stocks; patients submit vaccine requests including personal information; vaccination centers update storage capacity and vaccination progress; and carriers register their service conditions. Table I contains part of the data these participants must provide to the system.
In a subsequent stage (data usage), authorized participants retrieve and read the data provided earlier, either specific to single entities or aggregated, and receive decision support depending on their role in the process. For example, patients can retrieve their declarations at any time, carriers can be informed about the on-time delivery rate, and the medical hub requests automated support for vaccine supply distribution. Decisions related to the vaccine distribution include (i) the priority to be assigned to patients to administer the vaccine dose (which we will use as a running example throughout this paper), (ii) the restocking need for each vaccination center, (iii) and which carrier should handle each delivery.
In our example, several critical aspects need to be considered. First, all data is confidential and must be protected from unauthorized access. For example, revealing a single patient’s metadata constitutes a privacy breach. Second, data integrity and availability are critical. Any unauthorized modification could alter decisions or falsify patient records, while system failures or data unavailability could disrupt the decision-making process. Ensuring confidentiality, integrity, and availability while enabling a secure and verifiable automated decision-making is an open challenge motivating our research.
Table II presents the requirements identified from the motivating use case that underpins our approach.
Table II: Requirements and corresponding actions
Requirement Approach
R1 The stored information shall be tamper-proof, traceable, confidential, permanent, and available We store the encrypted data in a tamper-proof distributed file storage, and we use a notarization system to keep track of information
R2 Decision support shall be user-definable, trusted, reproducible, verifiable, and confidential We handle data within trusted hardware, encrypting them before being stored outside of it; data access and processing is established through customizable policies, turned into verifiable apps running in trusted hardware
R3 The decision support system must be secure, interoperable, available, replicable, and scalable We employ standard languages to describe authorization policies (expressing what each user can do and what resources they may access) and decision logic (the rules to follow in making decisions)
Ensuring that decision-critical information is consistently available, tamper-evident, and auditable is essential to both the efficiency of healthcare processes and trust in the system. The information source must be immutable and accessible, with traceability ensured to facilitate subsequent validations and verifications. We formalize these needs in Tab. II. Since the system manages sensitive and confidential healthcare data, its security is crucial to ensuring correct decisions, and prevent unauthorized access, data leaks, malicious attacks, and system unavailability. For example, disclosing patients’ metadata or their prioritization in receiving the vaccines would violate their privacy. We enclose these properties in Tab. II. Finally, our approach must be interoperable with existing data storage, access, and aggregation solutions while enabling secure and automated decision-making support for confidential data. We meet these aspects with Tab. II. Our solution aims to meet these requirements.
IV Approach
In this section, we present our approach named Secure Platform for Automated decision Rules via Trusted Applications (SPARTA). We first introduce the system components (Sect. IV-A), and then describe how they interact to manage information and enable secure decision-making over confidential data (Sect. IV-B).
IV-A Actors and basic components
[…]
Figure 1: An overview of the SPARTA approach with the software components, user roles, and main information flow
Figure 1 provides a bird’s eye view of the SPARTA approach. We identify four main actors, each responsible for a distinct functionality. The Policymaker defines the decision rules and specifies which functions are available, along with the participant attributes or attribute combinations required to invoke them. Considering the example in Sect. III, the Policymaker role can be fulfilled either by a regulatory body such as the Ministry of Health, or by the consortium directly involved in the decision process (upon consensus, as in [36]). The Data Provider provides the input data and contextual information used for decision-making. In our example, the Data Providers are the participants: the medical hub, patients, vaccination centers, and the carriers. The Certification Authority (CA) signs a Certificate Signing Request (CSR) generated by a Registration Authority (RA), thereby producing a valid certificate. The RA handles participant registration and verifies their identity and role. Once verification is complete, it generates the CSR. By signing the CSR, the CA attests actor metadata, authenticates participants, and certifies their roles and attributes, determining their eligibility to interact with the system and invoke specific functions. In our example, the role of the CA should be fulfilled by a globally recognized Certification Authority (e.g., DigiCert, Entrust, or GlobalSign), [^6] DigiCert: digicert.com; Entrust: entrust.com; GlobalSign: globalsign.com/en, accessed 2025-08-04 akin to what is offered by the National Health Service in the United Kingdom [37]. The RA role should be covered by a National Health Service Provider or a government eID system for patients; a Regional Department of Public Health or Healthcare Ministry for vaccination centers; and a Ministry of Transport or logistics regulator for carriers. Finally, the Decider receives decision support based on the available data. If permitted by the access policies, the Decider may also retrieve raw data and invoke aggregation functions. In our example, the Decider is the medical hub.
The interplay of five basic components realizes the architecture. (I) The Confidential Computing Unit (CCU) is SPARTA’s core component. We deploy the CCU inside a Trusted Execution Environment (TEE). This hardware-secured environment guarantees data confidentiality and ensures the integrity of both computation and code (see Sect. II), thereby meeting Tab. II. Notice that we make CCU replicable for the sake of information availability and system robustness (Tabs. II and II). The CCU deploys in a dedicated enclave within the TEE a separate Decision Trusted App (DeTApp) for each decision process. A DeTApp consists of the following. During SPARTA’s execution phase, it exposes (i) an invocable Decision Support Object (DeSObj) providing and controlling access to the Policymaker’s decision-making functions. The DeSObj, in turn, calls (ii) a Decision Function Object (DeFObj) realizing those functions, and a linked (iii) Decision Library (DeLib), a predefined library that offers utilities for DeSObjs such as data retrieval and decryption, policy checking, and certificate validation. The first two software objects are customizable and are built in SPARTA’s initialization phase by a DeSObj builder and a DeFObj builder, respectively, based on a policy descriptor and a decision model provided by the Policymaker. A DeTApp builder binds those objects together with the DeLib and deploys the DeTApp in the enclave.
(II) The Request Manager is a façade to the CCU. It can be abstracted as a queue that handles both requests to and responses from the CCU. It is designed to ensure the preservation of request order while guaranteeing that the CCU processes one request at a time to circumvent issues related to concurrent computation. Notice that the Request Manager is not within a secure environment like the TEE, so data must be encrypted whenever it is sent to it. The Request Manager handles calls from the two following interface components.
(III) The Admin Interface handles information exchange with CA and Policymaker. (IV) The User Interface is the means through which the Data Providers and the Decider interact with the CCU. Finally, SPARTA is paired with (V) the Data Manager, an external component providing two architectural buttresses: (i) The Data File Processor, storing the content of the exchanged data, and (ii) The Notarization System, saving and sealing the resource locators to those data. To meet Tab. II, we leverage two technologies for the realization of the above components. We employ IPNS over IPFS for the Data File Processor because IPNS guarantees permanence, immutability and content-addressing for stored files (see Sect. II). We designate a public blockchain platform as the Notarization System to store the IPNS names in a tamper-proof, available and traceable way. We remark that other solutions could be used such as STORJ777storj.io, accessed 2025-07-30 and distributed relational databases [38], respectively, in scenarios with less restrictive security requirements.
Next, we elucidate the interplay of the above components.
IV-B Workflow
The workflow of SPARTA mainly consists of two phases, as illustrated in Fig. 1: First, the system initialization (Sect. IV-B1) prepares the necessary data structures and programs, then the execution follows (Sect. IV-B2).
IV-B1 System initialization
The system initialization phase goes through four consecutive steps: (a) Certification(CERT), to let the attributes of participants be attested to; (b) Seed exchange(SDEX), to define a seed for encryption key generation; (c) Communication channel configuration(CONF), and (d) Specification of functions and policies(SPEC). We will detail them in the following.
Table III: Process participants’ and CCU types and attributes
Subject Participant Types Attributes
Central medical hub Data Provider, IdentificationNumber="IT-001-2000",
Decider Role="MedicalHub", Country="Italy"
Patient Data Provider IdentificationNumber="IT-001-3891",
Role="Patient", Country="Italy"
Vaccination center Data Provider IdentificationNumber="IT-001-5343",
Role="VaccinationCenter", Country="Italy"
Carrier Data Provider IdentificationNumber="IT-001-2435",
Role="Carrier", Country="Italy"
CCU Role="CCU"
Certification (CERT)
In this step, the CA provides evidence of the attributes held by actors using a certificate. Table III lists the types and attributes of the Data Providers involved in Sect. III and of the CCU component, too. For example, the CA attests that the medical hub who is both a Data Provider and a Decider, holds attributes specifying an IdentificationNumber, and that it is a MedicalHub operating in Italy. This certificate contains the Elliptic Curve Digital Signature Algorithm (ECDSA) [39] public key of the actor and their attributes. Issuing the certificate, the CA attests to this information for each user. In our approach, we implemented a CA that simulates the issuance of X.509 certificates, signing them with its ECDSA private key. Listing 1 shows an example of such a certificate. Each participant must check the validity of the counterpart’s certificate with the CA’s ECDSA public key before any interaction (as per Tab. II). Furthermore, the certificate has a unique identification number, a time validity, and the user’s ECDSA public key. Notably, the CCU must also obtain a valid certificate from the CA to let the users check whether they are interacting with the correct computation unit (meeting Tab. II).
Listing 1: Example of X.509 certificate issued by the CA
1X.509 certificate {
2 SerialNumber: 2000,
3 Subject: Central Medical Hub,
4 ParticipantTypes: [Data Provider, Decider],
5 Validity: 1 year,
6 Attributes:
7 IdentificationNumber="IT-001-2000",
8 Role="MedicalHub",
9 Country="Italy",
10 IDCheck=True
11 ECDSA public key: "MFkw[...]WXpA==" }
A note on the remote attestation (RA)
Before making any request, a process participant like the Data Provider, the Decider, or even other CCUs (henceforth, a client) performs the remote attestation to establish trust with the CCU deployed inside a TEE (as per Tab. II). This phase follows the principles defined in the RATS RFC standard [40], the foundation for several TEE attestation schemes, such as Intel EPID and AMD SEV-SNP. Remote attestation serves a dual objective: on one hand, provide participants with strong evidence that the data they are transmitting is sent to a Trusted Application running within a TEE; on the other hand, verify the identity and integrity of the DeSObj as an authentic CCU software entity, recognized as such by the CA (see the CERT step) within the context of the running process (in this case, the scenario described in Sect. III). We slightly modify the remote attestation, to meet all the features described. To perform it, the client (i) retrieves and checks the CCU’s certificate released by the CA to verify that the CCU is the intended one and not a malicious attacker. Then, (ii) the client retrieves and validates the TLS certificate issued by the CCU to establish a secure connection with it, and finally (iii) gets (via attested TLS channel) and verifies the attestation report of the CCU. The attestation report contains the so-called measurement of the Trusted App, which is defined as the hash value of the combination of its source code and data. This report is signed using the attestation private key associated with the CCU’s hardware. To verify it, the client (a) authenticates the CCU using the public key and decrypts the report, (b) compares the measurement found within the decrypted report against a predefined reference value associated with the source code of the CCU (if the decrypted measurement matches the predefined value, the CCU gains trust from the client), (c) verifies that the report data matches the server’s Transport Layer Security (TLS) certificate.
Seed exchange (SDEX)
Relying on a single CCU introduces significant vulnerabilities that could compromise Tabs. II and II. If this component becomes unavailable due to failure, downtime, or malicious attacks, the entire execution process could halt indefinitely, potentially making it impossible to recover. Additionally, if that CCU exclusively holds the cryptographic keys, all associated data could be permanently lost. We recall that the Request Manager is not within a TEE, so information exchanges with the system should be encrypted with a known key Sect. IV-A. To deal with these issues in SPARTA, the CCU creates and shares with other CCUs a seed to ensure that they all generate the necessary cryptographic keys deterministically (meeting Tab. II). To explain how we achieve this, let us consider a scenario with two CCUs: CCU1 and CCU2. Let us assume that CCU1 generates the seed and sends it to CCU2. CCU1 does so as follows. (1) It performs the remote attestation of CCU2. If the remote attestation (RA ) succeeds, (2) it generates a secure random string (the seed) and seals it with a sealing key known only to itself (see Sect. II); then, it sends the seed to CCU2 via the secure channel established during RA . As CCU2 receives the secret seed from CCU1, it performs the first two operations as well, namely it performs the remote attestation of CCU1, and, if the latter succeeds, it seals the seed with a sealing key known only to itself. At that point, the exchanged seed transmitted by CCU1 can be shared with an arbitrary number of other CCUs. We will describe how the seed is used to generate encryption keys next, and in Sect. IV-B2.
Communication channel configuration (CONF)
After the seed exchange, a CCU can generate a certificate for establishing secure communication channels to and from the Request Manager via TLS over TCP, so that confidential data is not leaked (as per Tab. II). More specifically, we slightly modify the TLS certificate, allowing it to contain, as an extra parameter, an Elliptic-Curve Diffie-Hellman (ECDH) [41] public key. The CCU loads the seed created in the SDEX step and uses it to generate a deterministic ECDH private/public key pair (using the X25519 curve [42]). Notice that with more than one CCU, since all of them share the same seed (see the SDEX step), they generate the same ECDH key pair. This ECDH key is used by the client to generate a shared secret before transmission to the Request Manager (see Sect. IV-B2). Since the CCU signs the TLS certificate with its ECDSA private key and the corresponding ECDSA public key is present inside the certificate, every participant can verify the signature with the CCU’s ECDSA public key. The ECDSA public key is certified by the CA (see the CERT step), so a participant can trust both the certificate and the CCU. At the end of this step, the CCU is ready to receive requests and provide responses through the Request Manager.
Specification of functions and policies (SPEC)
After the certification phase, the Policymaker declares the decision functions the system should realize in a decision model, and their access rules in a policy descriptor. In order to meet Tab. II, the Policymaker should be allowed to express those with standard, well-established, and customizable languages. Here we resort to DMN with FEEL, and ALFA, respectively (see Sect. II). An explanation of DMN, FEEL, and ALFA scripting languages transcends the scope of this paper. It suffices to explain their use here by example.
Table IV: Decision model for the patient prioritization of Sect. III
When And And And And And And And Then
Age PreExistingConditions CurrentMedications PreviousVaccinations FamilyMedicalHistory ConsentFormSigned meanAge sumAge
PatientPrioritization
WithAggr
number string string string string boolean number number string
$\geq$60 “Asthma”,“Diabetes” “Metformin”,“Albuterol” “COVID-19”,“Influenza” “Diabetes”,“Heart Disease” true $\geq$40 >250 “High”
[18..60[ “Asthma” “Metformin” “Influenza” - true <40 - “Medium”
<18 - “Lisinopril” “COVID-19” - true ]18..60[ - “Low”
- - - - - false - - “Ineligible”
Table IV shows an example of a DMN indicating how a decision is made for the prioritization of vaccine requests by patients. All input columns refer to data attributes in the structure retaining patient information. Among them, the meanAge and sumAge columns are aggregation functions. For the sake of readability, we separately show their encoding in FEEL in Tab. V. Considering the meanAge aggregation, e.g., the FEEL formula filters the data (based, among others, on the patients’ family medical history, age, and previous vaccinations), and then computes the mean age of the remaining entries. The DeFObj builder (see Sect. IV-A) transforms such a table into executable code.
Listing 2: Decision function object derived from Tab. IV
[…]
2 if inputs["Age"] >= 60 and (inputs["PreExistingConditions"] == "Asthma" or inputs["PreExistingConditions"] == "Diabetes") and (inputs["CurrentMedications"] == "Metformin" or inputs["CurrentMedications"] == "Albuterol") and (inputs["PreviousVaccinations"] == "COVID-19" or inputs["PreviousVaccinations"] == "Influenza") and (inputs["FamilyMedicalHistory"] == "Diabetes" or inputs["FamilyMedicalHistory"] == "Heart Disease") and inputs["ConsentFormSigned"] == true and inputs["MeanAge"] >= 40 and inputs["SumAge"] > 250 {return "High"}
3 if inputs["Age"] >= 18 and inputs["Age"] < 60 and inputs["PreExistingConditions"] == "Asthma" and inputs["CurrentMedications"] == "Metformin" and inputs["PreviousVaccinations"] == "Influenza" and inputs["ConsentFormSigned"] == true and inputs["MeanAge"] < 40 {return "Medium"}
4 if inputs["Age"] < 18 and inputs["CurrentMedications"] == "Lisinopril" and inputs["PreviousVaccinations"] == "COVID-19" and inputs["ConsentFormSigned"] == true and inputs["MeanAge"] > 18 and inputs["MeanAge"] < 60 {return "Low"}
5 if inputs["ConsentFormSigned"] == false {return "Ineligible"}
[…]
Listing 2 shows the result for our example. Notice that the client requesting this decision will not see the underlying data but only observe the result, because that code will run within the CCU in a TEE.
Listing 3: An example of policy descriptor in ALFA
1namespace VaccineDispatch {
2 import Oasis.Attributes.*
3 attribute Country{id="Country" category=environmentCat type=string}
[…]
7 apply denyUnlessPermit
8 rule accessDecision {
9 permit
[…]
Table V: FEEL formulas the functions in List. 3
Aggr. func. FEEL formula
filteredPatients:
Patient[ConsentFormSigned = true and
contains(FamilyMedicalHistory, "Heart Disease")
and (Age <= 38 or (Age > 40 and
contains(PreExistingConditions, "Diabetes"))) and
contains(PreviousVaccinations, "COVID-19")],
meanAge: mean(filteredPatients.Age)
filteredPatients:
Patient[ConsentFormSigned = true],
sumAge: sum(filteredPatients.Age)
[…]
1
Input: descriptorFile containing: funcName, the string representing the name of the function; accPol, the function’s access policy; datStNam, the indexing name of the input data structure; aggrFuncs, a (possibly empty) list of aggregation functions on the input data.
Output : The DeSObj script to be deployed inside the CCU and invocable by the process participants. The « symbol denotes the direct yield of a string output.
2
3begin
[…]
5  « $(\textit{cert},\textit{func},\textit{ipnsName})\leftarrowDeLib$.ParseDecisionReq(payload)
6  « $(\textit{certVal},\textit{attr})\leftarrow$ DeLib.CheckCertificate(cert)
7  « if certVal==true then
[…]
9     « if calVal==true then
[…]
11        « $\textit{aggrInputs}\leftarrow$ []
12        foreach aggrF in aggrFuncs do
[…]
14           « aggrInputs $\leftarrow$ Append(aggrInputs,aggrVar)
[…]
16        « return decision
17    « else throw exception “Access policy not satisfied”
18
19 « else throw exception “Invalid certificate”
20
21
Algorithm 1 Code generator for DeSObj instances
Listing 4 presents the DeSObj script to run inside the CCU produced by the DeSObj builder (Algorithm 1) using List. 3 as input.
Listing 4: A DeSObj script generated via Algorithm 1 with the policy descriptor in List. 3 as an input: funcName is assigned with PatientPrioritizationWithAggr, accPol with isMedicalHub, dataStruct with Patient, and aggrFuncs with an array consisting of meanAge and sumAge (Tab. V).
[…]
4 if certVal {
[…]
6 if calVal {
[…]
8 aggrInputs := make(map[string]interface{})
[…]
10 aggrInputs["MeanAge"] = MeanAge
[…]
12 aggrInputs["SumAge"] = SumAge
[…]
15 return "Access policy not satisfied" }
16 return "Invalid certificate" }
Notice that this script does not contain direct calls to the decision functions, or data retrieval. It delegates the actual execution to the DeLib instead. The code inside the DeLib is parametric and independent of the use case scenario. Only the DeSObj script is tailored to the specific decision process. Finally, the DeTApp builder deploys the above software objects within the TEE enclave as a DeTApp.
To sum up, at the end of this step a Decider can invoke a function like List. 4. The automated decision-making (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:decision) will be carried out by the evaluation of a DeFObj like List. 2 from within the enclave during the execution phase, which we discuss next.
IV-B2 Execution
The lilac box encloses the components involved in the execution phase of SPARTA Fig. 1. It encompasses a writing phase, and a decision making phase. In the writing stage (WS) , data is provided for decisions, meeting Tabs. II and II. It is comprised of two steps, respectively corresponding to the data provision and data usage stages in our motivating scenario (Sect. III).
Client encryption (WS-i)
This step basically establishes a secure asynchronous communication channel over an insecure channel, i.e., the Request Manager’s queue, between the client and the CCU. It pertains to the Data Provider when sending the input data to the CCU through the Request Manager. As specified in Sect. IV-A, the Request Manager is not trusted, so data must be encrypted before its transmission; otherwise, another participant or a malicious attacker can sniff data from the queue. To avoid this, the Data Provider, when performing the remote attestation against the CCU, retrieves the CCU’s ECDH public key from the TLS certificate (see the CONF step). Then, they generate an ECDH private/public key pair [41] using the X25519 curve [42] and use their own private ECDH key along the CCU’s ECDH public key to generate a shared secret thanks to the ECDH algorithm. From the shared secret, the Data Provider generates a symmetric key deterministically by using Secure Hash Algorithm (SHA) (in particular SHA-256 [43, 44]) and uses this key to encrypt the data before sending it to the CCU via the Request Manager using AES-GCM [45, 46]. Moreover, the Data Provider signs their own ECDH public key with their ECDSA private key (see the CERT step) and attaches the signed ECDH public key and the ECDSA public key in the payload. After reading the input data from the request queue, the CCU checks the CA certificate of the invoking Data Provider (using the ECDSA public key of the CA), then checks the ECDH signature with the ECDSA public key (of the Data Provider), and finally generates the shared secret thanks to its private ECDH key and the Data Provider’s public ECDH key with the ECDH algorithm. The shared secret will be equal to the one calculated by the Data Provider in the encryption phase. From that secret, it generates the symmetric key, and since this operation is deterministic, this key will also be equal to the one calculated by the Data Provider. At this point, the CCU decrypts the data sent by the Data Provider. The symmetric key exchange mechanism described above guarantees that the Request Manager, even if fully corrupted, cannot learn the key shared between the Data Provider and the CCU, thus ensuring that the Request Manager cannot read any message encrypted using a secure symmetric key encryption mechanism such as AES under the shared symmetric key.
CCU encryption (WS-ii)
The second writing phase pertains to the CCU and follows the Data Provider’s input data decryption. To securely store data in a public IPFS, the CCU generates a new unique symmetric key from a pre-determined private seed each time a new element must be saved to IPFS. Since these keys are deterministically computed from the CCU’s private seed, the CCU can easily recompute the very same key each time a particular piece of data must be decrypted for computing the final decisions. In practice, to compute this derived symmetric key, the CCU generates a random number and calculates the SHA-3 hash [43] of the seed shared during the SDEX step, and the random number. Finally, it uses this symmetric key to encrypt the data with AES-GCM (meeting Tab. II,Tab. II). The encrypted data is put on IPFS. The random number must be freshly generated for each record and must be the same as the one used in the decryption phase described next. The fact that each random number for each entry is written on IPFS in clear does not impact security since the SHA-3 operation on each random number and the seed produces a new symmetric key which is totally unrelated from the CCU’s private seed and the other keys. To meet Tab. II, the CCU stores all the necessary IPNS names in a smart contract on a public blockchain. It is noteworthy to say that since the CCU encrypts data also using a seed established during the SDEX step, malicious attackers who were not part of this phase cannot impersonate a CCU to access the data, as they cannot steal the seed to reconstruct the decryption key because it is encrypted with a sealing key (see Sect. II)—and they cannot be part of the SDEX procedure without a valid certificate form the CA.
Decision-Making Support (DECM)
This phase corresponds to the data usage stage described in Sect. III. When a Decider asks for decision support, they invoke the decision function of the DeSObj deployed inside the CCU as per Tab. II (e.g., the PatientPrioritizationWithAggrHandler in List. 4). Recall that every request is handled by the Request Manager. Before sending the request, the participant performs the remote attestation against the CCU (see the note on RA). If the attestation completes successfully, the participant finalizes the request. The payload of the request contains the Decider’s certificate released by the CA, the name of the function they are invoking (in this case, PatientPrioritizationWithAggr), and the IPNS key name (e.g., PatientsWAggrKey). As in List. 4, after parsing the participant request with the DeLib (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:payload), the CCU checks the received CA certificate of the requester invoking the DeLib (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:check-certificate). If it is valid, then the CCU extracts the attributes inside the certificate and checks whether they satisfy the function’s access policy linked to the invoked decision function thanks to the DeLib (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:check-callability). In our example, the access policy is isMedicalHub. If the attributes of the Decider are the ones of List. 1, the access policy of the invoked function is satisfied by the attributes the Decider bears, so they can invoke it. At this stage, the CCU retrieves and decrypts the data with the corresponding function inside the DeLib (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:decrypt-data) and performs the two aggregations functions evaluating the corresponding FEEL formulas invoking the DeLib (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:mean-aggregationList. 4, ln. LABEL:lst:FinalDecisionWithAggregation:sum-aggregation). Finally, it invokes the desired decision function (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:decision, note that the PatientPrioritizationWithAggr decision function corresponds to the List. 2). Finally, the CCU sends the result back to the Request Manager that returns it to the Decider (List. 4, ln. LABEL:lst:FinalDecisionWithAggregation:return).
V Formal Analysis
In this section, we formalize SPARTA as a protocol to discuss its security guarantees. In Sect. V-A we present the necessary tools for the discussion, namely the notation, the Random Oracle and finally the Authenticated Encryption. Then, in Sect. V-B we introduce the properties that we require for our protocol (see Sect. V-B1). We then present a thread model of our solution (see Sect. V-B2), and we finally conclude with the Theorem and its proof in Sect. V-B3.
V-A Useful Tools
(1) Notation.Let $\mathbb{N}$ be the set of all natural numbers. We denote the security parameter by $\lambda\in\mathbb{N}$. Every algorithm takes the security parameter $\lambda$ as input (in unary, i.e., $1^{\lambda}$). When an algorithm has multiple inputs, $1^{\lambda}$ is typically omitted. A function is called negligible if, for every positive polynomial and all sufficiently large security parameter $\lambda$, the function grows at most as the inverse of the polynomial in $\lambda$. (2) Random Oracle.In the Random Oracle Model (ROM), all the parties have access to a truly random hash function. In particular, when a value $v$ is given as an input from a party to the Random Oracle (RO), the latter samples a random answer $r$, stores the pair $(v,r)$, and outputs $r$ to the party. If the RO is queried on the same value $v$ multiple times, the same answer $r$ is output. It is heuristically shown that cryptographic hash functions such as $\mathsf{SHA\text{-}256}$ behave as a random oracle. (3) Authenticated Encryption.A Symmetric Key Encryption scheme with key space $\mathcal{K}$ is composed of the following algorithms: (i) $\mathsf{Gen}(1^{\lambda})$: The randomized key generation algorithm outputs a key $k$ in the key space $\mathcal{K}$. (ii) $\mathsf{Enc}(k,m)$: The randomized encryption algorithm, on input the key $k$ and a message $m$, outputs a ciphertext $c$. (iii) $\mathsf{Dec}(k,c)$: The decryption algorithm, on input the key $k$ and a ciphertext $c$, outputs a message $m$, or $\bot$ if decryption fails. An authenticated encryption scheme should satisfy the following properties: (i) Correctness:For each message $m$ and each $k\in\mathcal{K}$, $\Pr[\mathsf{Dec}(k,\mathsf{Enc}(k,m))=m]=1$. (ii) Unforgeability:The probability that a PPT adversary $\mathcal{A}$ wins the following experiment (i.e., outputs 1) is negligible: (a) Obtain $k$ from $\mathsf{Gen}(1^{\lambda})$. (b) The adversary, after making at most polynomially many queries to the encryption oracle $\mathsf{Enc}(k,\cdot)$, outputs a ciphertext $c$. (c) Let $m=\mathsf{Dec}(k,c)$. Output 1 if $m\neq\bot$ and $c$ was not queried to the encryption oracle by $\mathcal{A}$. Output 0 otherwise. (iii) CCA-Security:The probability that a PPT adversary $\mathcal{A}$ wins the following experiment (i.e., outputs 1) is negligibly close to $1/2$: (a) Obtain $k$ from $\mathsf{Gen}(1^{\lambda})$. (b) The adversary, with oracle access to $\mathsf{Enc}(k,\cdot)$ and $\mathsf{Dec}(k,\cdot)$, outputs messages $m_{0},m_{1}$. (c) Choose a uniform bit $b\in\{0,1\}$, compute ciphertext $c=\mathsf{Enc}(k,m_{b})$, and send $c$ to $\mathcal{A}$. (d) The adversary, with oracle access to $\mathsf{Enc}(k,\cdot)$ and restricted oracle access to $\mathsf{Dec}(k,\cdot)$ (excluding $c$), outputs a bit $b^{\prime}$. (e) The experiment outputs 11 if $b^{\prime}=b$, else 0.
V-B Formal Analysis
V-B1 Desired Properties
We require the following properties from our protocol:
(1) Correctness.For any function $f$ and any set of messages $M$ such that $f(M)=y$, if $M$ is provided to the CCU for function $f$ by the Data Provider obtaining ciphertexts $C$, and subsequently, ciphertexts $C$ are provided by the Decider back to the CCU obtaining output $y^{\prime}$, then $y^{\prime}=f(M)=y$. (2) Data Privacy.A PPT adversary $\mathcal{A}$ wins the following experiment with probability negligibly close to $1/2$: (i) $\mathcal{A}$chooses a function $f$. $\mathcal{A}$ has oracle access to the CCU acting as Data Provider (sending arbitrary sets of messages $M$ and receiving corresponding ciphertexts $C$) and oracle access to the CCU acting as Decider (sending arbitrary sets of ciphertexts $C$ and receiving back $f(M)$ if $M$ correctly decrypts from $C$). Then, $\mathcal{A}$ chooses two sets of messages $M_{0},M_{1}$ such that $f(M^{\prime}_{0})=f(M^{\prime}_{1})$ where $M^{\prime}_{b}$ can be any combination of messages in $M_{b}$. (ii) Choose a random bit $b$, produce ciphertexts $C_{b}$ of messages $M_{b}$, and send $C_{b}$ to $\mathcal{A}$. (iii) $\mathcal{A}$has oracle access to the CCU as a Data Provider to send arbitrary sets of messages $M$, and obtain the corresponding ciphertexts $C$ back. It has further oracle access to the CCU as a Decider by sending arbitrary ciphertexts $C$ and obtaining back $f(M)$ if $M$ correctly decrypts $C$. Finally, outputs a bit $b^{\prime}$. (iv) If $b^{\prime}=b$, output 1; otherwise, output 0. (3) Data Integrity.A PPT adversary $\mathcal{A}$ wins the following experiment with negligible probability: (i) $\mathcal{A}$, acting as Data Provider, has oracle access to the CCU submitting arbitrary message sets $M$ and receiving ciphertexts $C$. $\mathcal{A}$ outputs ciphertexts $C^{*}$. (ii) Decrypting ciphertexts $C^{*}$ yields messages $M^{*}$. Output 1 if no message in $M^{*}$ equals $\bot$ and no subset of $M^{*}$ was previously queried; else, output 0. (4) Forward/Backward Secrecy/Integrity.If all ciphertext keys at time $t$ are leaked, ciphertexts encrypted at any different time $t^{\prime}$ remain secure, and their underlying messages cannot be decrypted or forged.
V-B2 Threat Model
We assume the Data Provider and Certification Authority are honest, the Decider is semi-honest, [^8] We could assume a dishonest Decider by adding signatures on the decision and data produced by the CCU, but this is unnecessary in our setting. and secure authenticated channels can be established between authorized parties (e.g., CA, Decider) and the CCU. We model the CCU as a fully trusted third party computing decisions through functions $\{f_{j}\}$ using a hardwired secret seed $s$.
V-B3 The Protocol
Let $(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec})$ be an authenticated encryption scheme with key space $\mathcal{K}$ and $H:\{0,1\}^{*}\rightarrow\mathcal{K}$ be a cryptographic hash function modeled as a random oracle. The protocol is as follows:
(1) Setup:The CCU is initialized with functions $\{f_{j}\}$ and a uniform seed $s\in\{0,1\}^{\lambda}$. (2) Data Storage: (i) The Data Provider establishes a secure channel with the CCU, chooses message set $M$, and sends it to the CCU. (ii) The CCU generates a unique ID $T_{i}$, key $k_{i}=H(s,T_{i})$, ciphertexts $c_{i}=\mathsf{Enc}(k_{i},m_{i})$ for $m_{i}\in M$, and stores ciphertexts publicly. (3) Decision: (i) Decider establishes a secure channel with the CCU and sends ciphertext set $C$ and function index $j$. (ii) CCUtakes the unique ID $T_{i}$, keys $k_{i}=H(s,T_{i})$, decrypts ciphertexts to messages $m_{i}$, and outputs decision $f_{j}(\{m_{i}\})$.
We are now ready to prove the following informal theorem.
Theorem 1.
Assuming that $(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec})$ is an Authenticated Encryption scheme and $H$ is modeled as a Random Oracle, the protocol described above satisfies Correctness, Data Privacy, Data Integrity, and Forward/Backward Secrecy and Integrity.
Proof.
Recall that the CCU is modeled as a fully trusted third party. Hence, the fixed functions $\{f_{j}\}$ cannot be modified, internal states remain secret, and outputs are always correctly computed with respect to these functions.
Correctness is ensured by the correctness property of the authenticated encryption scheme $(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec})$ and by the fact that the CCU’s functions are fixed and unmodifiable.
Data Privacy is guaranteed since the internal CCU state (i.e., the seed $s$) remains secret and the derived encryption keys $k_{i}=H(s,T_{i})$ appear indistinguishable from random due to the Random Oracle assumption. Thus, Data Privacy reduces directly to the CCA-security of $(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec})$.
Data Integrity follows from the Unforgeability property of the authenticated encryption scheme $(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec})$, combined with the trusted execution guarantees provided by the CCU.
Forward/Backward Secrecy and Integrity result directly from the random-oracle-based key derivation: given a leaked key $k=H(s,T)$, the one-way property of the Random Oracle prevents recovery of the seed $s$. Therefore, keys corresponding to ciphertexts encrypted at other timestamps $T^{\prime}\neq T$ remain computationally infeasible to derive. ∎
With correctness and key security properties formally proven, we now focus on describing the SPARTA proof-of-concept we implemented and its evaluation.
VI Implementation and evaluation
We implemented SPARTA in Go, using the EGo framework to deploy trusted applications within an Intel SGX TEE. [^9] EGo: edgeless.systems/products/ego; Intel SGX: sgx101.gitbook.io/sgx101. Accessed 2025-06-17 We realized the Request Manager as a Redis queue. To evaluate FEEL filtering statements and access policies we resort to govaluate. For the evaluation of aggregation functions we rely on Expr. [^10] Redis queue: github.com/rq/rq; govaluate: github.com/Knetic/govaluate; Expr: github.com/expr-lang/expr. Accessed 2025-06-17 Our tests were conducted on a TEE hosted by an Intel Xeon Gold 5415+ CPU.
The objective of our experimental evaluation is to demonstrate the feasibility of our solution and assess its performance in multiple application scenarios. We gauge not only runtime but also computation memory in light of the limits imposed by TEEs (see Sect. II). In particular, we aim to answer the following core questions: 1) Does SPARTA scale efficiently with increasing input data size and decision table complexity? 2) Can SPARTA operate end-to-end on realistic input data and policy configurations across multiple domains? 3) What is the computational overhead introduced by encryption and the use of a TEE? To answer these questions and test distinct features of our system, we used the following three datasets.
VAX (synthetic vaccine dataset).
We generated this dataset based on our vaccine campaign use case (see Sect. III), incorporating realistic access control policies and data aggregation requirements relevant to public health scenarios. This synthetic dataset is essential to assess the practical feasibility of SPARTA and, due to the numerous data fields, to measure the memory savings achieved through our aggregation mechanism described in the WS step.
MED (real-world medical dataset).
This dataset originates from the February 2021 Decision Management Community challenge Determine Medical Service Coverage and contains over 16 00016\,000 healthcare-related records. Its large size makes it suitable for testing the scalability of SPARTA under realistic conditions, including measuring execution time with large input and evaluating performance when varying the complexity and size of the decision tables. The nature of the data (medical services and patient records) aligns well with scenarios requiring strong data confidentiality, allowing us to test encryption strategies and directly compare SPARTA’s performance inside a TEE versus a traditional server execution.
CARD (real-world financial dataset).
This dataset comes from the October 2021 Decision Management Community challenge Card Approval Decision. Its decision logic includes diverse input conditions, such as relational operators (e.g., >>, $\geq$, $\leq$, <<) and interval-based conditions. The presence of those predicates is essential for testing the flexibility of SPARTA in handling complex decision-making scenarios. Additionally, using a dataset from a financial context allows us to validate that our approach is adaptable to other sensitive domains.
Next, we illustrate our experimental results.
VI-1 Running example tests
We assess the practical feasibility of SPARTA and quantify the memory savings achieved by the data aggregation mechanism (see the WS phase) using the synthetic VAX dataset, which simulates our motivating scenario of a nationwide vaccine campaign. Following the setup in Sect. III, we generated realistic input data for patients, vaccination centers, and carriers (Tab. I), aligning each record with the decision logic defined in the corresponding DMN tables (see Tab. IV for our running example, and the appendix for the other two decision functions). Each rule in the decision logic is matched by at least one input, and we manually verify the correctness of the computed decisions. To ensure reproducibility, all data generators, inputs, and outputs are publicly available in our repository.111111anonymous.4open.science/r/SPARTA
[…]
Figure 2: Execution time with increasing records and columns
VI-2 Determine Medical Service Coverage tests
We evaluate SPARTA’s scalability and execution performance when dealing with large datasets by using MED (16 36916\,369 rules, $7$ input columns, and $3$ output columns corresponding to the final decisions). We measure the performance impact under varied conditions, specifically by increasing (a) the number of records and columns, and (b) the number of rules. In addition to the core scalability tests, we use the same dataset to measure: (c) the performance overhead of our aggregation function, in terms of both execution time and (d) memory usage, (e) the encryption overhead required to provide strong confidentiality guarantees, and (f) the overhead of using a TEE compared to a standard, non-secure server execution. To execute the tests, we randomly select $300$ rules (an arbitrary number in line with the name of our approach, SPARTA) out of 16 36916\,369.
(a) Execution time with increasing number of records and columns
Figure 3 reports the execution time of the decision function (List. 2) under increasing workload. In Fig. 3(a), we fix the number of columns in the DMN to 771414, and $28$, and progressively increase the number of input records from $1000$ to 16 00016\,000. In Fig. 3(b), we fix the number of records to 1000100080008000, and 16 00016\,000, and vary the number of columns from 1 to 28. In both configurations, we observe a linear increase in execution time as the number of records and columns grows. The linear scalability shows that our solution scaled predictably with increasing data size and complexity. Such a characteristic is essential for real-world deployments where decision tables and input data may grow significantly.
Figure 3: Decision execution time
(b) Execution time with increasing number of rules
To assess the impact of the complexity of the decision on SPARTA’s runtime, we perform a test by incrementally increasing the number of rules involved. Figure 5 shows the result with a fixed number of records (16 00016\,000) and columns ($28$) while varying the number of rules from $100$ up to $1500$. The observed trend is linear, thus confirming a good scalability of our system also with regard to the number of decision rules.
(c) Execution time with function aggregation results as additional inputs
As described in the SPEC step, the Policymaker may define aggregation functions to be computed prior to the decision process, with the resulting values used as additional inputs. To test the overhead introduced by this feature, we extend the original $7$ columns by adding up to $21$ aggregation functions for a total of $28$ columns as input. From the eighth column onward, each column corresponds to one the aggregation output. Figure 5 shows the result with fixed 16 00016\,000 records. The figure shows that the decryption and aggregation time introduce a small overhead to the computation. The total time is comparable to the decision time. This result entails that the decryption and aggregation time are features that can be included in the system adding a small time overhead.
[…]
Figure 4: Execution time with increasing number of rules
[…]
Figure 5: Execution time with prior aggregation
(d) Memory usage
We measure the memory usage within the TEE (the enclave memory) during full execution runs. Figure 6 shows the memory Random Access Memory (RAM) usage for two different scenario: Figure 6(a) presents the results with prior aggregation, while Fig. 6(b) without. Each run begins with the initialization of the CCU and ends when the decision process terminates. The two images demonstrate a reasonable memory footprint throughout the computation with an expected spike in correspondence with data decryption.
Figure 6: RAM usage over time with and without prior aggregation
(e) Encryption overhead
To assess the overhead, we conducted experiments that account not only for the decision-making process itself but also for the data decryption phase. This step is essential in our system, as data must be decrypted after retrieval from IPFS and before decision execution (see the DECM step). We design two test configurations to evaluate the impact of different encryption strategies: light encryption, which serves as a baseline for comparison, and heavy encryption. Recalling the WS-ii step, in the heavy configuration, the CCU generates a different random number for each record, so it encrypts each such record with a different symmetric key (as per Tabs. II and II). For the light encryption variant, we modified the process described in WS-ii to use a single symmetric key (namely, we used the same random number for generating the symmetric key) for all entries. Indeed, this setup is strictly for experimental purposes. In practice, such an approach is discouraged, as compromising the single key would allow an attacker to decrypt the entire dataset rather than a single record. Figure 7 shows the execution time of the two encryption solutions. Figure 7(a) displays the execution time of light and heavy encryption strategies by varying the number of records with a fixed amount of columns ($28$), while in Fig. 7(b) we increase the number of columns with a fixed amount of users (16 00016\,000). The results highlight that the heavy encryption approach incurs a modest but reasonable overhead ($\approx\qty$80ms at most) compared to the light variant. However, this performance trade-off is justified because per-record encryption ensures significantly higher data confidentiality and resilience against key compromise.
Figure 7: Comparison between light and heavy encryption
(f) TEE vs Server
To quantify the overhead introduced by using both TEE and strong cryptographic guarantees, we compare our decision support system running inside the TEE with a standard server-based execution, inspired by [47, 48]. Specifically, we modify our system to run outside the TEE but on the same machine. Figure 8 shows the comparison result: Figure 8(a) presents the execution time with increasing number of records and fixed columns, while in Fig. 8(b) we vary the number of columns with a fixed number of records (16 00016\,000). The results shows that the additional overhead is limited and primarily due to the decryption phase required by the TEE. In contrast, the server execution assume direct access to plaintext data within a trusted environment. Overall, our system provides strong security guarantees, including confidentiality for both data and decision logic, at a minimal performance cost and without compromising efficiency.
Figure 8: Execution time comparison between server and TEE
VI-3 Card Approval tests
Finally, to evaluate SPARTA’s ability to support varied and complex decision logic, we use the CARD dataset, which models a financial decision-making scenario. This dataset features a DMN table containing $3$ input columns and $13$ rules, including relational and interval-based conditions, operators often used in real-world access control and policy systems. As no input data is provided with the original challenge, we synthetically generate inputs that ensure full rule coverage. To assess performance at scale, we replicate these inputs up to 100 000100\,000 records.
Figure 9: Execution time with VAX, MED, and CARD datasets
Figure 9 summarizes execution times across all evaluated datasets. Notably, the MED dataset incurs higher decision evaluation time due to its larger rule set (see Sect. VI-2). However, decryption time remains consistent, underscoring SPARTA’s ability to maintain stable performance across domains with varying logic complexity and input scale.
VII Limitations and Future Work
Our solution presents some limitations. Firstly, our current design involves a single TEE. Even though SPARTA allows for the replication of CCUs, a load balancing strategy is not yet implemented. Such an extension would require non-trivial communication and session management adjustments, but would allow for a large-scale adoption.
The memory limits of a TEE (see Sect. II) presents a scalability challenge when it comes to make decisions over a large volume of records. Solutions to this problem can involve integrating Oblivious RAM (ORAM) techniques [49, 50, 51, 52] or the execution of decisions in batched segments that can circumvent those memory constraints.
We developed our solution for Intel SGX TEEs. Other solutions exist (e.g., AMD SEV, Arm CCA) and studying their employment is an interesting endeavor. Also, we will consider the integration of Intel Trust Domain Extensions (TDX). Our implementation does not support the entire FEEL language to write the filtering expressions and the aggregation functions. In particular, we currently back only one operation over the filtered data, and support only the standard aggregation functions (e.g., meansummaxmin). To further increase the expressive power of the decision making processes, we envision the adoption of TEE-based smart contracts for automated decision-making [53, 54, 55, 56, 57].
Finally, we conducted experiments with public benchmarks showing good performance and scalability of our approach with low overhead. Nevertheless, an on-field use of our system in real-world settings is paramount for validation purposes.
VIII Related Work
In recent years, numerous approaches have been proposed in the field of decision-making support systems. Haarmann et al. [14] propose executing DMN decision models on the Ethereum blockchain. However, blockchain-based approaches inherently rely on public ledgers, making them unsuitable for confidential data applications. In contrast, our TEE-based solution enables automated decision-making support securely and privately while maintaining the trust, transparency, and audibility required in multi-party processes. Moreover, in our solution, functions’ access policies ensure that only authorized participants can interact with the system, thus receiving decision-making support or confidential data. Zhang et al. [58] introduce a privacy-preserving decision-making system over blockchain for secure voting. In this paper, the authors proposed a low complexity Distributed Batch Key Generation and then designed new Two-Stage Voting scheme to reduce voting effort. This solution solves the privacy issues that arise when using blockchain technology, but it is tailored for decision-making voting only. On the contrary, SPARTA allows for user-definable decision logic. Other solutions, such as [13] and [12], allow for privacy-preserving clinical decisions for automated classification. These solutions resort to homomorphic encryption to maintain data confidentiality, resulting in significant performance overhead due to encryption. They resort to a Naïve Bayes classifier and a Gaussian kernel-based SVM as the decision support system. In contrast, our solution uses AES-GCM encryption, introducing low overhead to the overall performance and allowing for specifying the decision logic for the decision-making support. Liu et al. [59] present HPCS, a Hybrid Privacy-Preserving Clinical Decision Support System used for health status monitoring and disease prediction. Here, a server securely monitors patients’ conditions in real-time. The new symptoms can be sent to the cloud server running a neural network for prediction in a privacy-preserving way. Different from SPARTA, the authors use the Paillier Cryptosystem with Threshold Decryption to guarantee data privacy, adding high overhead on the overall procedure. Moreover, HPCS limits the decision logic to the sole predictions, while we allow for user-definable logic. Xie et al. [60] present TEBDS, a data-sharing system combining TEEs with blockchain. Their design leverages a TEE for off-chain storage and a consortium blockchain for access control in IoT contexts, with a centralized key management center (KMC). In contrast, our approach relies on a public blockchain only for storing IPNS names pointing to IPFS content-addressed data while delegating all operations to the TEE. This design eliminates the need for a centralized KMC and enables confidential decision-making over encrypted data.
IX Conclusion
In this paper, we presented SPARTA, an approach that enables automated decision-making support over confidential data in multi-party settings while preserving the secrecy of the information. SPARTA leverages a TEE-based architecture to securely execute decision functions and access policies expressed using DMN and ALFA, respectively. By executing these software objects within trusted hardware, SPARTA ensures confidentiality, integrity, and verifiability without exposing sensitive data. We discuss its security guarantees with a threat model analysis and formal proof. We also test our prototype, measuring time, memory, performance overhead, and scalability with synthetic and real-world datasets.